OpenStudy (anonymous):

The function g(x,y) = x^4+y^3 has a critical point at (0,0). What sort of critical point is it?

OpenStudy (amistre64):

its either a min, a max, or a saddle that is a min of one and a max of the other id assume

OpenStudy (amistre64):

gx = 4x^3 gxx = 12x^2 gxy = 0 gy = 3y^2 gyy = 3y gyx = 0

OpenStudy (amistre64):



OpenStudy (anonymous):

The second derivative test fails but i there is another method the has something to do with completing the square

OpenStudy (amistre64):

if we graph the projections into the zx and zy planes we can see that its a min

OpenStudy (amistre64):

err, something close to a min :)

OpenStudy (amistre64):